OpenStudy (anonymous):

Identify the correct slope and y intercept of the equation 6x - 3y = 12. slope = 2; y intercept at (0, -4) slope = -2; y intercept at (0, 4) slope =1/2 ; y intercept at (0, 15) slope =1/2 ; y intercept at (0, 9)

OpenStudy (anonymous):

6x - 3y = 12 Put in slope intercept form y = mx + b Subtract 6x from both sides -3y = -6x + 12 Divide both sides by -3 y = 2x - 4 Therefor the slope is 2 and the y-int is -4 Answer A slope = 2; y intercept at (0, -4)

OpenStudy (anonymous):

ok thank you for explainig it

OpenStudy (anonymous):

Np :)